# Gradient-Guided Annealing for Domain Generalization

### v6

@@ 4.4 Sensitivity Analysis
Regarding the magnitude of weight perturbations during the application of GGA, we found that the optimal value was $\rho=5e-5$. As illustrated in Figure 4, a larger magnitude of perturbation led to decreased model performance. Intuitively, the application of larger noise to the model parameters leads to sets that are not close to the solution, making it increasingly difficult for the model to converge. On the other hand, smaller perturbations seem to have little to no effect on training, as the search is limited to spaces near the current parameters, which is why the model performance falls back close to that of ERM. With regards to the stage of training during which GGA will be applied, we found that the models yielded better performance when the search was initialized in earlier stages.We hypothesize that applying perturbations near the end of training displaces the model from a local optimum, requiring additional iterations to converge.
@@ G.2 Complexity Analysis
H Full Experimental Results
@@
In this section, we show detailed results of Table 2 in the main text. The results marked by $\dagger$,$\ddagger$ , ‡ are copied from Gulrajani and Lopez-Paz [31] and Wang et al. [81], respectively. Standard errors for the baseline methods are reported from three trials, if available from past literature. In green and red, we highlight the performance boost and decrease of applying GGA on top of each algorithm respectively, averaged over three trials. In addition, we also present detailed results for the DomainNet benchmark, without however including results for the combination of GGA with the baseline algorithms, due to computational restrictions. We also include experiments for the ColoredMNIST and RotatedMNIST datasets, where we reproduced the results for all baselines and report the average results over 5 runs. The below tables are better read in color.

### v7

@@ 4.4 Sensitivity Analysis
Regarding the magnitude of weight perturbations during the application of GGA, we found that the optimal value was $\rho=1e-5$. As illustrated in Figure 4, a larger magnitude of perturbation led to decreased model performance. Intuitively, the application of larger noise to the model parameters leads to sets that are not close to the solution, making it increasingly difficult for the model to converge. On the other hand, smaller perturbations seem to have little to no effect on training, as the search is limited to spaces near the current parameters, which is why the model performance falls back close to that of ERM. With regards to the stage of training during which GGA will be applied, we found that the models yielded better performance when the search was initialized in earlier stages.We hypothesize that applying perturbations near the end of training displaces the model from a local optimum, requiring additional iterations to converge.
@@ G.2 Complexity Analysis
H An Alternative Method for Gradient-Guided Annealing: GGA-L
In this section we propose we propose an alternative method to GGA, which integrates noise directly into the gradient update step rather than modifying weights separately. The key idea is to inject controlled noise during vanilla optimization, guiding updates toward regions with more stable gradient similarity across domains. Normally, in SGD, parameters update as: $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\cdot\nabla\mathcal{L}_{\mathcal{B}}(\mathcal{B};\bm{\theta}_{t})$ for a sample batch $\mathcal{B}$. Instead of perturbing model weights manually for specific training steps and checking gradient similarity iteratively, as in GGA, we propose injecting dynamic noise based on domain gradient similarity directly into the update step, as follows:
$\bm{\theta_{t+1}}=\bm{\theta_{t}}-\eta\cdot(\nabla\mathcal{L}_{\mathcal{B}}(\mathcal{B};\bm{\theta_{t}})+\alpha\cdot\bm{\xi})$ (7)
where $\bm{\xi}\sim\mathcal{U}(0,1)$ is noise drawn from the Uniform distribution and
$\alpha=\gamma\cdot(1-\frac{2}{\lvert\mathcal{S}\rvert(\lvert\mathcal{S}\rvert-1)}\sum_{\begin{subarray}{c}i=1,j=1\\ i\neq j\end{subarray}}^{\lvert\mathcal{S}\rvert}\left(\frac{\bm{g}^{T}_{i}\cdot\bm{g}_{j}}{\|\bm{g}_{i}\|\|\bm{g}_{j}\|}\right))$ (8)
is a dynamic scaling factor depending on the average gradient similarity between domains and $\gamma$ is a hyperparameter controlling the noise intensity. The noise term $\alpha$ acts as an implicit exploration, pushing the optimizer toward parameter sets where the gradient similarity is naturally higher. When gradients from different domains are already similar (i.e., high cosine similarity), $\alpha$ remains small, whereas if the gradients diverge, $\alpha$ increases, adding more noise to encourage exploration in the parameter space. As the above proposed method is inspired by Stochastic Gradient Langevin Dynamics (SGLD) [84], we call it GGA-L.
Following the same experimental protocol and hyperparameters as in the main paper, we show the results on PACS, VLCS, OfficeHome and DomainNet in Table 4. We set the $\gamma$ hyperparameter to $31e-$ 3 for all experiments expect DomainNet where we set it to $41e-$ 4 and apply the noisy update rule throughout training. For batch sizes we use 32 for PACS, VLCS and DomainNet and 48 for OfficeHome and TerraIncognita. As apparent from the results, GGA-L is able to perform similarly to GGA but adds a considerably lower computational cost as the gradients for each domain are only calculated once per batch. GGA-L is presented in Algorithm 2.
Algorithm 2 Implementing GGA-L in training
1:Pretrained DNN $h$ with parameters $\bm{\theta}_{0}$, training dataset with domains $\mathcal{S}=\{D_{i}\}_{1}^{\lvert\mathcal{S}\rvert}$. Loss gradients $\bm{g}$. Learning rate $\eta$, noise $\xi$ drawn from the Uniform distribution $\mathcal{U}(0,1)$, noise scaling factor $\alpha$, hyperparameter $\gamma$ controlling the noise intensity. Total number of training iterations $n$.
2:
3:for $t\leftarrow 1$ to $n$ do
5:       Compute the mini-batch loss at starting params:
6:          $\mathcal{L}_{B}\leftarrow\mathcal{L}(\mathcal{B};\bm{\theta}_{t})$
7:       #GGA-L/
8:       Calculate minimum domain grad pair sim:
9:       $\text{sim}\leftarrow\displaystyle\frac{2}{\lvert\mathcal{S}\rvert(\lvert\mathcal{S}\rvert-1)}\sum\limits_{\begin{subarray}{c}i=1,\,j=1\\ i\neq j\end{subarray}}^{\lvert\mathcal{S}\rvert}\left(\frac{\bm{g}^{T}_{i}\cdot\bm{g}_{j}}{\|\bm{g}_{i}\|\|\bm{g}_{j}\|}\right)$
10:       Calculate noise scaling factor $\alpha$
11:          $\alpha=\gamma\cdot(1-\text{sim})$
12:       #Update weights:
13:       $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\cdot\nabla(\mathcal{L}_{\mathcal{B}}(\mathcal{B};\bm{\theta}_{t})+\alpha\cdot\bm{\xi})$
14:       $t=t+1$
15:end for
Table 4: Comparison of GGA and GGA-L with the ERM baseline.The top out-of-domain accuracies on five domain generalization benchmarks averaged over three trials, are presented. We report the average over PACS, VLCS, OfficeHome and TerraInc in the “Avg.’ column, while the average over all 5 datasets in the final one.
Algorithm PACS VLCS OfficeHome TerraInc Avg. DomainNet Total Avg.
ERM 85.5±$\pm 0.2$ 77.3±$\pm 0.4$ 66.5±$\pm 0.3$ 46.1±$\pm 1.8$ 68.9 43.8±$\pm 0.3$ 63.9
GGA 86.4±$\pm 0.5$ 78.7±$\pm 0.8$ 67.0±$\pm 0.3$ 48.5±$\pm 1.1$ 70.2 44.5±$\pm 0.2$ 65.0
GGA-L 86.5±$\pm 0.7$ 78.4±$\pm 0.9$ 66.5±$\pm 0.4$ 49.8±$\pm 1.8$ 70.3 44.5±$\pm 0.2$ 65.1
I Full Experimental Results
@@
In this section, we show detailed results of Table 2 in the main text. The results marked by $\dagger$,$\ddagger$ , ‡ are copied from Gulrajani and Lopez-Paz [31] and Wang et al. [81], respectively. Standard errors for the baseline methods are reported from three trials, if available from past literature. In green and red, we highlight the performance boost and decrease of applying GGA on top of each algorithm respectively, averaged over three trials. In addition, we also present detailed results for the DomainNet benchmark, without however including results for the combination of GGA with the baseline algorithms, due to computational restrictions. We also include experiments for the ColoredMNIST and RotatedMNIST datasets and report the average results over 5 runs.